A 60g bullet is fired at a target that presents an average resistive force to the passage of the bullet of 20kN per metre thickness of the material from which is constructed. If the bullet strikes the target at 600ms-1, what does the minimum thickness of the target have to be to stop the passage of the bullet completely?


1
Expert's answer
2022-02-21T12:06:02-0500

Given:

"m=0.060\:\rm kg"

"F=20000\:\rm N"

"v=600\:\rm m/s"


The energy work theorem says

"0-\frac{mv^2}{2}=-Fd"

Hence, the thickness of the target

"d=\frac{mv^2}{2F}=\frac{0.060*600^2}{2*20000}=0.54\:\rm m=54\: cm"